Exercise 1. If x is a real number, prove that there are integers p, q ∈ Z such that p < x < q.


1
Expert's answer
2021-01-19T01:58:49-0500

Let q=p+1, then


"p<x\le q"


"p<x\le p+1"

If x=p+1, then we will take

"q=p+i\times 1"

where i is any natural nomber.

Then we will have any I for will be true next:


"p<x<p+i\times 1"

For example,